Given a normal distribution with mean equal to 30 and standard deviation of 6, find the normal curve are to the left of x=22


1
Expert's answer
2021-05-17T17:50:01-0400

Z multiplier =(X-mu)/sigma

=(22-30)/6=-1.33

Negative value it's self means area belong to left of x=22

Probability corresponding to z=-1.33 is 0.098